OpenStudy (anonymous):

Find the slope of the line that contains the following points. A(5, 6), B(10, 8)

OpenStudy (ryan123345):



OpenStudy (anonymous):

slope of line is 2/5 hence equation of line is y=mx+c for c any point can be used 8=10*1/5+c c=6 hence y=1/5*x+6 and 5y=x+30

OpenStudy (anonymous):

14/15

OpenStudy (ryan123345):

@autumnpoteet , what is 8-6 and 10-5 ?

OpenStudy (anonymous):

2/5?

OpenStudy (ryan123345):

Good Job :)